# Bayesian Optimization of Robustness Measures Using Randomized GP-UCB-based Algorithms under Input Uncertainty

### v1

@@ B.2 Theoretical Analysis in Uncontrollable Settings when $\Omega$ is Finite
Under the assumption in Theorem B.3, the following holds:
@@
where $\hat{t}$ is given by (4.3), and functions $q_{1}(\cdot)$, $h_{i}(\cdot)$ and each coefficient are given in Theorem B.3. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
@@ B.3 Extension to Uncontrollable Settings when $\Omega$ is Continuous
Let $\Omega$ be a continuous set. In this case, if $\mathcal{X}$ is finite or continuous, the only difference between the proposed method and Algorithm 3 or 4 is whether or not ${\bm{w}}_{t}$ is sampled from $P^{\ast}$. We give the pseudo-code for the case where $\Omega$ is continuous in the uncontrollable setting in Algorithm 6. Note that $\mathcal{X}$ in Algorithm 6 includes both the finite and continuous cases.
Algorithm 6 RRGP-UCB for robustness measures in the uncontrollable setting when $\Omega$ is continuous
@@
   Generate $\xi_{t}$ from the chi-squared distribution with two degrees of freedom
@@
   Estimate $\hat{\bm{x}}^{\dagger}_{t}$ by $\hat{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu^{\dagger}_{t-1}({\bm{x}},{\bm{w}}))$
   Select the next evaluation point $\bm{x}_{t}$ by (A.3)
@@
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at the point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding the observed data
@@ B.4 Theoretical Analysis in Uncontrollable Settings when $\Omega$ is Continuous
Assume that Regularity assumption, Assumption A.3, A.4, A.5, B.2 and B.3, and (A.1) hold. Suppose that $\mathcal{X}$ and $\Omega$ are finite and continuous, respectively. Let $\tau^{\dagger}_{t}=\lceil b_{2}d_{2}rt^{2}(\sqrt{\log(a_{2}d_{2}|\mathcal{X}|)}+\sqrt{\pi}/2)\rceil$, and let $\Omega_{t}$ be a set of discretization for $\Omega$ with each coordinate equally divided into $\tau^{\dagger}_{t}$. Suppose that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow $P^{\ast}$, and $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Define $\kappa^{(5)}_{t}=\lceil b_{2}d_{2}rt^{2}(\sqrt{\log(a_{2}d_{2}|\mathcal{X}|)}+\sqrt{\pi}/2)\rceil^{d_{2}}|\mathcal{X}|$ and $\beta_{t}=2\log\kappa^{(5)}_{t}+\xi_{t}$. Let $q_{2}(a)$ and $q^{\dagger}(a)=\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumption A.4 and A.5, respectively. For the sequence $\iota_{1},\ldots,\iota_{t}$ satisfying Assumption B.3, define $\varphi_{t}=\iota_{1}+\cdots+\iota_{t}$. Then, under the uncontrollable setting, the following holds if Algorithm 6 is performed:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})}$, $C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, $p^{-1}_{min,t}$ is given in Assumption B.3, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\beta_{1},\ldots,\beta_{t}$. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
@@
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for other measures.
@@
Under the assumption in Theorem B.5, for $\hat{t}$ defined by (A.4), the following holds:
@@
In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
@@
Assume that Regularity assumption, Assumption A.5, A.6, A.7, B.2 and B.3, and (A.1) hold. Suppose that $\mathcal{X}$ and $\Omega$ are continuous. Let $\tilde{\tau}_{t}=\lceil b_{3}drt^{2}(\sqrt{\log(a_{3}d)}+\sqrt{\pi}/2)\rceil$, and let $\mathcal{X}_{t}\times\Omega_{t}$ be a set of discretization for $\mathcal{X}\times\Omega$ with each coordinate equally divided into $\tilde{\tau}_{t}$. Suppose that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow $P^{\ast}$, and $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Define $\kappa^{(5)}_{t}=(1+\tilde{\tau}^{d_{1}}_{t})\tilde{\tau}^{d_{2}}_{t}$ and $\beta_{t}=2\log\kappa^{(5)}_{t}+\xi_{t}$. Let $q_{3}(a)$ and $q^{\dagger}(a)=\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumption A.7 and A.5, respectively. For the sequence $\iota_{1},\ldots,\iota_{t}$ satisfying Assumption B.3, let $\varphi_{t}=\iota_{1}+\cdots+\iota_{t}$. Then, under the uncontrollable setting, the following holds if Algorithm 6 is performed:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, $p^{-1}_{min,t}$ is given in Assumption B.3, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\beta_{1},\ldots,\beta_{t}$. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
@@
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for other measures.
@@
Under the assumption in Theorem B.7, for $\hat{t}$ defined by (A.4), the following holds:
@@
In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
@@ C Proofs
The value $\max_{{\bm{x}}\in\mathcal{X}}{\rm l}_{t-1}({\bm{x}})$ is a constant that does not depend on ${\bm{x}}$. This implies that $\tilde{\bm{x}}_{t}={\bm{x}}^{({\rm u})}_{t}$. Therefore, if ${\bm{x}}_{t}=\tilde{\bm{x}}_{t}$, then ${\rm u}_{t-1}({\bm{x}}^{({\rm u})}_{t})={\rm u}_{t-1}({\bm{x}}_{t})$. On the other hand, if ${\bm{x}}_{t}=\check{\bm{x}}_{t}$, that is,
$2\beta^{1/2}_{t}\sigma_{t-1}(\tilde{\bm{x}}_{t})={\rm u}_{t-1}(\tilde{\bm{x}}_{t})-{\rm l}_{t-1}(\tilde{\bm{x}}_{t})\leq{\rm u}_{t-1}(\check{\bm{x}}_{t})-{\rm l}_{t-1}(\check{\bm{x}}_{t})=2\beta^{1/2}_{t}\sigma_{t-1}(\check{\bm{x}}_{t}),$
then $\beta^{1/2}_{t}\sigma_{t-1}(\tilde{\bm{x}}_{t})\leq\beta^{1/2}_{t}\sigma_{t-1}(\check{\bm{x}}_{t})$. From the definition of $\check{\bm{x}}_{t}$, $\mu_{t-1}(\tilde{\bm{x}}_{t})\leq\mu_{t-1}(\check{\bm{x}}_{t})$. Since $\beta^{1/2}_{t}\sigma_{t-1}(\tilde{\bm{x}}_{t})\leq\beta^{1/2}_{t}\sigma_{t-1}(\check{\bm{x}}_{t})$, the inequality $\mu_{t-1}(\tilde{\bm{x}}_{t})+\beta^{1/2}_{t}\sigma_{t-1}(\tilde{\bm{x}}_{t})\leq\mu_{t-1}(\check{\bm{x}}_{t})+\beta^{1/2}_{t}\sigma_{t-1}(\check{\bm{x}}_{t})$ holds. This implies that ${\rm u}_{t-1}(\tilde{\bm{x}}_{t})\leq{\rm u}_{t-1}(\check{\bm{x}}_{t})$. Finally, since $\tilde{\bm{x}}_{t}={\bm{x}}^{({\rm u})}_{t}$, ${\rm u}_{t-1}({\bm{x}}^{({\rm u})}_{t})\leq{\rm u}_{t-1}(\check{\bm{x}}_{t})$ and ${\rm u}_{t-1}({\bm{x}}^{({\rm u})}_{t})={\rm u}_{t-1}(\check{\bm{x}}_{t})$ hold. ∎
@@
Hence, for the function $f({\bm{x}},{\bm{w}})$ with respect to ${\bm{w}}$, $f({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})$ holds. In addition, from the theorem’s assumption, since ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ satisfy (3.2), the following holds:
@@
where the third inequality is derived by $\mu_{t-1}({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})$, the definition of $\hat{\bm{x}}_{t}$, and $\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}})\leq\max_{{\bm{x}}\in\mathcal{X}}\rho(\mu_{t-1}({\bm{x}},{\bm{w}}))=\rho(\mu_{t-1}(\hat{\bm{x}}_{t},{\bm{w}}))$. Moreover, from Assumption 4.1, there exists a function
@@
$\mathbb{E}_{t-1}[F(\hat{\bm{x}}_{i})]=\rho(\mu_{t-1}(\hat{\bm{x}}_{i},{\bm{w}}))\leq\rho(\mu_{t-1}(\hat{\bm{x}}_{t},{\bm{w}}))=\mathbb{E}_{t-1}[F(\hat{\bm{x}}_{t})].$
@@
C.5 Equality between $q(a)$ in Table 4 and $q_{1}(a)$, $q_{2}(a)$, and $q_{3}(a)$
@@
For any ${\bm{x}},{\bm{x}}^{\prime}$, $f({\bm{x}},{\bm{w}})$ and $f^{\prime}({\bm{x}}^{\prime},{\bm{w}})$, we show
$|\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))|\leq q\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right).$
Since ${\bm{x}},{\bm{x}}^{\prime},f({\bm{x}},{\bm{w}})$ and $f^{\prime}({\bm{x}}^{\prime},{\bm{w}})$ are arbitrary, we can assume that $|\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))|=\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))$ without loss of generality.
@@
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))$ $=\mathbb{E}_{{\bm{w}}}[f({\bm{x}},{\bm{w}})]-\mathbb{E}_{{\bm{w}}}[f^{\prime}({\bm{x}}^{\prime},{\bm{w}})]$
@@
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))=\inf_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})-\inf_{{\bm{w}}\in\Omega}f^{\prime}({\bm{x}}^{\prime},{\bm{w}})\leq\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|.$
@@
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))=\sup_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})-\sup_{{\bm{w}}\in\Omega}f^{\prime}({\bm{x}}^{\prime},{\bm{w}})\leq\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|.$
@@
$\rho(f({\bm{x}},{\bm{w}}))=\inf\{b\in\mathbb{R}\mid\alpha\leq\mathbb{P}_{{\bm{w}}}(f({\bm{x}},{\bm{w}})\leq b)\}$ $\leq\inf\{b\in\mathbb{R}\mid\alpha\leq\mathbb{P}_{{\bm{w}}}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}})+k\leq b)\}$
@@
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))\leq k=\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|.$
@@
$\rho(f({\bm{x}},{\bm{w}}))=\frac{1}{\alpha}\int_{0}^{\alpha}v_{f}({\bm{x}};\alpha^{\prime}){\rm d}\alpha^{\prime}.$

### v2

@@ B.2 Theoretical Analysis in Uncontrollable Settings when $\Omega$ is Finite
Under the assumptions of Theorem B.3, the following holds:
@@
where $\hat{t}$ is given by (4.4), and functions $q_{1}(\cdot)$, $h_{i}(\cdot)$ and all coefficients are as defined in Theorem B.3. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following holds:
@@ B.3 Extension to Uncontrollable Settings when $\Omega$ is Continuous
Let $\Omega$ be a continuous set. In this case, if $\mathcal{X}$ is finite or continuous, the only difference between the proposed method and Algorithm 4 or 5 is whether or not ${\bm{w}}_{t}$ is sampled from $P^{\ast}$. We give the pseudo-code for the case where $\Omega$ is continuous in the uncontrollable setting in Algorithm 7. Note that $\mathcal{X}$ in Algorithm 7 includes both the finite and continuous cases.
Algorithm 7 RRGP-UCB for robustness measures in the uncontrollable setting when $\Omega$ is continuous.
@@
   Generate $\xi_{t}$ from chi-squared distribution with two degrees of freedom
@@
   Estimate $\hat{\bm{x}}^{\dagger}_{t}$ by $\hat{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho({\mu^{\dagger}_{t-1}({\bm{x}},\cdot)})$
   Select next evaluation point $\bm{x}_{t}$ by (A.3)
@@
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding observed data
@@ B.4 Theoretical Analysis in Uncontrollable Settings when $\Omega$ is Continuous
Assume that the regularity assumption, Assumptions A.3, A.4, A.5, B.2 and B.3, and (A.1) hold. Suppose that $\mathcal{X}$ and $\Omega$ are finite and continuous, respectively. Let $\tau^{\dagger}_{t}=\lceil b_{2}d_{2}rt^{2}(\sqrt{\log(a_{2}d_{2}|\mathcal{X}|)}+\sqrt{\pi}/2)\rceil$, and let $\Omega_{t}$ be a set of discretization for $\Omega$ with each coordinate equally divided into $\tau^{\dagger}_{t}$. Suppose that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow $P^{\ast}$, and $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Define $\kappa^{(5)}_{t}=\lceil b_{2}d_{2}rt^{2}(\sqrt{\log(a_{2}d_{2}|\mathcal{X}|)}+\sqrt{\pi}/2)\rceil^{d_{2}}|\mathcal{X}|$ and $\beta_{t}=2\log\kappa^{(5)}_{t}+\xi_{t}$. Let $q_{2}(a)$ and $q^{\dagger}(a)=\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumptions A.4 and A.5, respectively. For the sequence $\iota_{1},\ldots,\iota_{t}$ satisfying Assumption B.3, define $\varphi_{t}=\iota_{1}+\cdots+\iota_{t}$. Then, under the uncontrollable setting, if Algorithm 7 is performed, the following holds:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})}$, $C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, $p^{-1}_{min,t}$ is given in Assumption B.3, and the expectation is taken over all sources of randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\beta_{1},\ldots,\beta_{t}$. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following holds:
@@
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for the other measures.
@@
Under the assumptions of Theorem B.5, for $\hat{t}$ defined by (A.4), the following holds:
@@
In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following holds:
@@
Assume that the regularity assumption, Assumptions A.5, A.6, A.7, B.2 and B.3, and (A.1) hold. Suppose that $\mathcal{X}$ and $\Omega$ are continuous. Let $\tilde{\tau}_{t}=\lceil b_{3}drt^{2}(\sqrt{\log(a_{3}d)}+\sqrt{\pi}/2)\rceil$, and let $\mathcal{X}_{t}\times\Omega_{t}$ be a set of discretization for $\mathcal{X}\times\Omega$ with each coordinate equally divided into $\tilde{\tau}_{t}$. Suppose that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow $P^{\ast}$, and $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Define $\kappa^{(5)}_{t}=(1+\tilde{\tau}^{d_{1}}_{t})\tilde{\tau}^{d_{2}}_{t}$ and $\beta_{t}=2\log\kappa^{(5)}_{t}+\xi_{t}$. Let $q_{3}(a)$ and $q^{\dagger}(a)=\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumptions A.7 and A.5, respectively. For the sequence $\iota_{1},\ldots,\iota_{t}$ satisfying Assumption B.3, let $\varphi_{t}=\iota_{1}+\cdots+\iota_{t}$. Then, under the uncontrollable setting, if Algorithm 7 is performed, the following holds:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, $p^{-1}_{min,t}$ is given in Assumption B.3, and the expectation is taken over all sources of randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\beta_{1},\ldots,\beta_{t}$. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following holds:
@@
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for the other measures.
@@
Under the assumptions of Theorem B.7, for $\hat{t}$ defined by (A.4), the following holds:
@@
In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following holds:
@@ C Proofs
The value $\max_{{\bm{x}}\in\mathcal{X}}{\rm l}_{t-1}({\bm{x}})$ is a constant that does not depend on ${\bm{x}}$. This implies that ${\tilde{\bm{x}}^{(f)}_{t}}={\bm{x}}^{({\rm u})}_{t}$. Therefore, if ${\bm{x}}^{{(f)}}_{t}={\tilde{\bm{x}}^{(f)}_{t}}$, then ${\rm u}_{t-1}({\bm{x}}^{({\rm u})}_{t})={\rm u}_{t-1}({\bm{x}}^{{(f)}}_{t})$. On the other hand, if ${\bm{x}}^{{(f)}}_{t}={\hat{\bm{x}}^{(f)}_{t}}$, that is,
$2\beta^{1/2}_{t}\sigma_{t-1}({\tilde{\bm{x}}^{(f)}_{t}})={\rm u}_{t-1}({\tilde{\bm{x}}^{(f)}_{t}})-{\rm l}_{t-1}({\tilde{\bm{x}}^{(f)}_{t}})\leq{\rm u}_{t-1}({\hat{\bm{x}}^{(f)}_{t}})-{\rm l}_{t-1}({\hat{\bm{x}}^{(f)}_{t}})=2\beta^{1/2}_{t}\sigma_{t-1}({\hat{\bm{x}}^{(f)}_{t}}),$
then $\beta^{1/2}_{t}\sigma_{t-1}({\tilde{\bm{x}}^{(f)}_{t}})\leq\beta^{1/2}_{t}\sigma_{t-1}({\hat{\bm{x}}^{(f)}_{t}})$. From the definition of ${\hat{\bm{x}}^{(f)}_{t}}$, $\mu_{t-1}({\tilde{\bm{x}}^{(f)}_{t}})\leq\mu_{t-1}({\hat{\bm{x}}^{(f)}_{t}})$. Since $\beta^{1/2}_{t}\sigma_{t-1}({\tilde{\bm{x}}^{(f)}_{t}})\leq\beta^{1/2}_{t}\sigma_{t-1}({\hat{\bm{x}}^{(f)}_{t}})$, the inequality $\mu_{t-1}({\tilde{\bm{x}}^{(f)}_{t}})+\beta^{1/2}_{t}\sigma_{t-1}({\tilde{\bm{x}}^{(f)}_{t}})\leq\mu_{t-1}({\hat{\bm{x}}^{(f)}_{t}})+\beta^{1/2}_{t}\sigma_{t-1}({\hat{\bm{x}}^{(f)}_{t}})$ holds. This implies that ${\rm u}_{t-1}({\tilde{\bm{x}}^{(f)}_{t}})\leq{\rm u}_{t-1}({\hat{\bm{x}}^{(f)}_{t}})$. Finally, since ${\tilde{\bm{x}}^{(f)}_{t}}={\bm{x}}^{({\rm u})}_{t}$, ${\rm u}_{t-1}({\bm{x}}^{({\rm u})}_{t})\leq{\rm u}_{t-1}({\hat{\bm{x}}^{(f)}_{t}})$ and ${\rm u}_{t-1}({\bm{x}}^{({\rm u})}_{t})={\rm u}_{t-1}({\hat{\bm{x}}^{(f)}_{t}})$ hold. ∎
@@
Hence, ${f({\bm{x}},\cdot)}\in G_{t-1}({\bm{x}})$ holds. In addition, from the theorem’s assumption, since ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ satisfy (3.2), the following holds:
@@
where the third inequality is derived by ${\mu_{t-1}({\bm{x}},\cdot)}\in G_{t-1}({\bm{x}})$, the definition of $\hat{\bm{x}}_{t}$, and $\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}})\leq\max_{{\bm{x}}\in\mathcal{X}}\rho({\mu_{t-1}({\bm{x}},\cdot)})=\rho({\mu_{t-1}(\hat{\bm{x}}_{t},\cdot)})$. Moreover, from Assumption 4.1, there exists a function
@@
$\mathbb{E}_{t-1}[F(\hat{\bm{x}}_{i})]=\rho({\mu_{t-1}(\hat{\bm{x}}_{i},\cdot)})\leq\rho({\mu_{t-1}(\hat{\bm{x}}_{t},\cdot)})=\mathbb{E}_{t-1}[F(\hat{\bm{x}}_{t})].$
@@
C.5 Equality between $q^{(m)}(a)$ in Table 4 in Inatsu et al., 2024a and $q_{1}(a)$, $q_{2}(a)$, and $q_{3}(a)$
@@
For any ${\bm{x}},{\bm{x}}^{\prime}$, $f({\bm{x}},{\bm{w}})$ and $f^{\prime}({\bm{x}}^{\prime},{\bm{w}})$, we derive $q(a)$ satisfying
$|\rho({f({\bm{x}},\cdot)})-\rho({f^{\prime}({\bm{x}}^{\prime},\cdot)})|\leq q\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right).$
Since ${\bm{x}},{\bm{x}}^{\prime},f({\bm{x}},{\bm{w}})$ and $f^{\prime}({\bm{x}}^{\prime},{\bm{w}})$ are arbitrary, we can assume that $|\rho({f({\bm{x}},\cdot)})-\rho({f^{\prime}({\bm{x}}^{\prime},\cdot)})|=\rho({f({\bm{x}},\cdot)})-\rho({f^{\prime}({\bm{x}}^{\prime},\cdot)})$ without loss of generality. For the distributionally robust, monotone Lipschitz map, and weighted sum measures, we can use $q^{(m)}(a)$ in Table 4 in Inatsu et al., 2024a by introducing additional assumptions about monotonicity, concavity, and taking zero at point zero.
@@
$\rho({f({\bm{x}},\cdot)})-\rho({f^{\prime}({\bm{x}}^{\prime},\cdot)})$ $=\mathbb{E}_{{\bm{w}}}[f({\bm{x}},{\bm{w}})]-\mathbb{E}_{{\bm{w}}}[f^{\prime}({\bm{x}}^{\prime},{\bm{w}})]$
@@
$\rho({f({\bm{x}},\cdot)})-\rho({f^{\prime}({\bm{x}}^{\prime},\cdot)})=\inf_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})-\inf_{{\bm{w}}\in\Omega}f^{\prime}({\bm{x}}^{\prime},{\bm{w}})\leq\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|.$
This implies that $q(a)=a$.
@@
$\rho({f({\bm{x}},\cdot)})-\rho({f^{\prime}({\bm{x}}^{\prime},\cdot)})=\sup_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})-\sup_{{\bm{w}}\in\Omega}f^{\prime}({\bm{x}}^{\prime},{\bm{w}})\leq\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|.$
This implies that $q(a)=a$.
@@
$v_{f}({\bm{x}};\alpha)\equiv\rho({f({\bm{x}},\cdot)})=\inf\{b\in\mathbb{R}\mid\alpha\leq\mathbb{P}_{{\bm{w}}}(f({\bm{x}},{\bm{w}})\leq b)\}$ $\leq\inf\{b\in\mathbb{R}\mid\alpha\leq\mathbb{P}_{{\bm{w}}}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}})+k\leq b)\}$
@@
$\rho({f({\bm{x}},\cdot)})-\rho({f^{\prime}({\bm{x}}^{\prime},\cdot)})\leq k=\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|.$
This implies that $q(a)=a$.
@@
$\rho({f({\bm{x}},\cdot)})=\frac{1}{\alpha}\int_{0}^{\alpha}v_{f}({\bm{x}};\alpha^{\prime}){\rm d}\alpha^{\prime}.$